3/4-3/5x=2/5x+2/4
I'm not so good at fractions...

Answer by checkley71(8403)

3/4-3/5X=2/5X+2/4 OR -3/5X-2/5X=2/4-3/4 OR -5/5X=-1/4 OR -1/X=-1/4 OR X=4